OpenStudy (anonymous):

Carin has 4 albums of stickers. Each album has 30 scented stickers and 16 puffy stickers. How many stickers does Carin have in all?

OpenStudy (misty1212):

\[4\times (30+16)\] will do it

OpenStudy (misty1212):

puffy stickers?

OpenStudy (anonymous):

I don't even know haha

OpenStudy (anonymous):

Maybe ones with like velvet or like very thick and soft stickers

OpenStudy (anonymous):

Keira bought 7 containers of gems. Each container has 24 blue gems and 32 yellow gems. Which expression shows how many more yellow gems than blue gems are in the containers? A. 7 × 24 B. 7 × 32 C. (7 × 32) – (7 × 24) D. (7 × 24) + (7 × 32)

OpenStudy (anonymous):

Diana packed 9 first-aid kits. Each kit had 24 bandages and 16 antiseptic wipes. How many more bandages than antiseptic wipes did Diana pack?

OpenStudy (misty1212):

\[9\times 24-9\times 16\] is one way

OpenStudy (misty1212):

another is \[9\times (24-16)\]

OpenStudy (misty1212):

but the simplest way is \(9\times 8\)

OpenStudy (anonymous):

okay last one Four cargo ships transported automobiles between ports. Each ship transported 146 cars and 65 trucks. How many more cars than trucks were transported?

OpenStudy (misty1212):

\[4(146-65)\]

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=4%28146-65%29

OpenStudy (anonymous):

okay thank you SO much for you're help

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]